# Dynamically Adaptive Reasoning via LLM-Guided MCTS for Efficient and Context-Aware KGQA

### v3

@@ Introduction
While LLM-Guided Expansion effectively narrows the search space by selecting semantically relevant relations, it does not guarantee that all expanded paths are correct or meaningful in the broader reasoning context. As the search progresses, path semantics evolve dynamically, and early promising paths may later become irrelevant or misleading. To address this, Context-Aware Path Evaluation integrates a lightweight Transformer-based path scorer into the simulation phase of MCTS. This scorer leverages cross-attention to jointly encode the question and the current reasoning path, allowing for adaptive evaluation that captures evolving semantics. By assigning scores to simulated paths based on question-path alignment, this module enables more accurate path ranking throughout the search process.
Context-Aware Path Evaluator.
@@
Specifically, in the Simulation phase, we evaluate the quality of each candidate path constructed during MCTS rollouts. Given a question $q$ and a candidate relation path $p_{r}=(r_{1},r_{2},\ldots,r_{l})$, where $p_{r}$ is formed by sequentially selecting relations during the expansion steps, we first encode both the question and the path using a pre-trained LLM. Let $\mathbf{z}_{q}\in\mathbb{R}^{d}$ denote the embedding of the question and $\mathbf{z}_{r_{i}}\in\mathbb{R}^{d}$ denote the embedding of relation $r_{i}$. To capture the sequential structure of relation paths, we incorporate a learnable position encoding $\mathbf{e}_{i}^{\text{pos}}$ for each relation $r_{i}$. The final input sequence is constructed by combining each relation embedding $\mathbf{z}_{r_{i}}$ with positional encoding and feeding it into a Transformer encoder:
$\mathbf{E}_{p_{r}}=\mathrm{Transformer}([\mathbf{z}_{r_{1}}+\mathbf{e}_{1}^{\text{pos}},\ldots,\mathbf{z}_{r_{l}}+\mathbf{e}_{l}^{\text{pos}}]),$
@@
where $\mathbf{e}_{i}^{\text{pos}}=\mathbf{E}^{\text{pos}}[i]$ denotes the relative position encoding for the $i$-th hop in the path, drawn from a trainable embedding matrix $\mathbf{E}^{\text{pos}}\in\mathbb{R}^{L\times d}$, with $L$ as the maximum path length and $d$ as the embedding dimension. To further incorporate question-specific information, we apply a cross-attention mechanism, allowing the encoded path representation $\mathbf{e}_{p_{r}}$ to attend to the question embedding $\mathbf{z}_{q}$:
$\mathbf{H}=\mathbf{E}_{p_{r}}+\mathrm{CrossAttn}(\mathbf{E}_{p_{r}},\mathbf{z}_{q}),$
@@
$with\quad\text{CrossAttn}(\mathbf{E}_{p_{r}},\mathbf{z}_{q})=\text{softmax}({\mathbf{E}_{p_{r}}\cdot\mathbf{z}_{q}^{T}}/{\sqrt{d_{k}}})\cdot\mathbf{z}_{q}.$
@@
We then employ attention pooling over relation representation $\mathbf{H}$ to obtain the hidden states of relation path:
$\mathbf{s}_{p_{r}}=\sum_{i=1}^{l}\alpha_{i}\textbf{h}_{i},\quad\alpha=\mathrm{Softmax}(\text{MLP}(\mathbf{H})),$
@@
where $\mathbf{h}_{i}$ denotes the hidden state of the $i$-th relation and $\alpha_{i}$ is its learned attention weight. This pooling mechanism enables the model to selectively emphasize informative steps along the reasoning path. Finally, the pooled path representation $\mathbf{s}_{p_{r}}$ is concatenated with the question embedding $\mathbf{z}_{q}$, and the combined vector is fed into a multi-layer perceptron to compute the plausibility score of the question-path pair:
$S(q,p_{r})=\mathrm{MLP}([\mathbf{s}_{p_{r}};\mathbf{z}_{q}]).$ (3)
@@
This context-aware evaluation model dynamically scores partial reasoning paths by jointly considering the question and the relation sequence, offering accurate and context-sensitive guidance to the MCTS search process.
Pre-training of Evaluator.
@@
To train the context-aware path evaluation model, we construct supervision signals by generating positive and negative relation paths from local subgraphs. A path is labeled positive if it connects the head entity to a correct answer entity within a predefined hop limit. Negative paths are drawn from two sources: hard negatives that end near but do not reach the answer, and random negatives obtained via walks that avoid answer entities entirely. Each training instance is a triplet $(q,p^{+},p^{-})$, and sequences are zero-padded with attention masks for efficient batch training.
@@
The model computes a plausibility score $S(q,p)$ for each question-path pair and is optimized using the Pair-wise Ranking loss to encourage higher scores for positive paths:
@@
where $\sigma(\cdot)$ is the sigmoid function. This training strategy equips the evaluator with the ability to distinguish plausible reasoning paths, thereby improving the guidance signal during MCTS-based inference.
Path-based Dynamic Refinement
@@
To construct supervision signals for fine-tuning, we dynamically sample pseudo-path pairs ($\hat{p}’_{i}$, $\hat{p}’_{j}$) from the set of explored paths during MCTS. Instead of relying on the evaluator’s predictions, we assign pseudo-labels based on empirically grounded values derived from the search process. Specifically, for entity $e_{i}$ along a reasoning path $p_{r}$, we define its search value as: $w_{e_{i}}=\frac{w_{p_{r}}}{n_{e_{i}}}$, where $w_{p_{r}}$ is the cumulative reward from all rollouts passing through $p_{r}$, and $n_{e_{i}}$ is the visite count of entity $e_{i}$. Given a pair of paths, we assign pseudo-labels based on their relative values:
$\left(\hat{p}^{+},\,\hat{p}^{-}\right)=\begin{cases}(p’_{i},\,p’_{j}),&\text{if }w_{e_{i}}>w_{e_{j}},\\ (p’_{j},\,p’_{i}),&\text{otherwise}.\end{cases}$ (6)
The path evaluator is then fine-tuned using the PR loss in Eq. (4), encouraging higher scores for more promising paths.
Reasoning Process
The overall reasoning process is summarized in Appendix A. The framework begins by initializing the path evaluation model to distinguish between plausible and implausible reasoning paths derived from the knowledge graph, establishing a strong foundation for downstream search. During the dynamic MCTS process, the algorithm iteratively performs selection, expansion, simulation, and backpropagation. In the expansion step, an LLM-based planner adaptively selects the top-$k$ relations most relevant to the question, effectively steering the search toward semantically meaningful paths. The path evaluation model informs the simulation phase by prioritizing trajectories that are more likely to yield correct answers. To enable continual adaptation, pseudo-path pairs obtained during search are periodically used to refine the evaluator. Finally, entities reached by high-scoring reasoning paths are aggregated to construct the answer set.
Experiments
Experimental Settings
Datasets. To evaluate the effectiveness of DAMR, we conduct experiments on two widely used KGQA benchmarks: WebQSP (Talmor and Berant 2018) and CWQ (Yih et al. 2016). Following prior work (Sun et al. 2023; Liu et al. 2025), we uniformly sample 1,000 questions from the test sets of both datasets to evaluate the performance. More details about datasets are provided in Appendix C.
@@
Semantic Parsing
KV-Mem 46.7 34.5 18.4 15.7
EmbedKGQA 66.6 - 45.9 -
QGG 73.0 73.8 36.9 37.4
NSM 68.7 62.8 47.6 42.4
TransferNet 71.4 - 48.6 -
KGT5 56.1 - 36.5 -
Retrieval
GraftNet 66.4 60.4 36.8 32.7
PullNet 68.1 - 45.9 -
SR+NSM 68.9 64.1 50.2 47.1
SR+NSM+E2E 69.5 64.1 49.3 46.3
LLMs
Flan-T5-xl 31.0 - 14.7 -
Alpaca-7B 51.8 - 27.4 -
Llama3-8B 30.3 25.7 30.5 27.8
Qwen2.5-7B 28.4 23.7 25.9 24.1
ChatGPT 66.8 - 39.9 -
ChatGPT+CoT 75.6 - 48.9 -
LLMs+KGs
UniKGQA 77.2 72.2 51.2 49.0
DECAF 82.1 78.8 70.4 -
KD-CoT 68.6 52.5 55.7 -
Nutrea 77.4 72.7 53.6 49.5
ToG 81.9 76.0 68.5 60.2
RoG 80.8 70.8 57.8 56.2
KAPING 72.4 65.1 53.4 50.3
ReasoningLM 78.5 71.0 69.0 64.0
FiDeLis 84.3 78.3 71.5 64.3
GNN-RAG 80.8 70.8 57.8 56.2
DoG 65.4 55.6 41.0 46.4
DualR 81.5 71.6 65.3 62.1
DP 87.5 81.4 75.8 69.4
RwT 87.0 79.7 72.4 66.7
@@
Baselines. We compare DAMR with a comprehensive set of baselines. These baselines include: the semantic parsing methods, e.g., KV-Mem (Miller et al. 2016), EmbedKGQA (Saxena, Tripathi, and Talukdar 2020), QGG (Lan and Jiang 2020), NSM (He et al. 2021), TransferNet (Shi et al. 2021), and KGT5 (Saxena, Kochsiek, and Gemulla 2022); the retrieval-based methods, e.g., GraftNet (Sun et al. 2018), PullNet (Sun, Bedrax-Weiss, and Cohen 2019), SR+NSM (Zhang et al. 2022), and SR+NSM+E2E (Zhang et al. 2022); the general LLMs, including Flan-T5-xl (Chung et al. 2024), Alpaca-7B (Taori et al. 2023), Llama3-8B (Dubey et al. 2024), Qwen2.5-7B (Team 2024), ChatGPT (Schulman et al. 2022), and ChatGPT+CoT (Wei et al. 2022); and recent LLMs with KG methods, including UniKGQA (Jiang et al. 2022), DECAF (Yu et al. 2022), KD-CoT (Wang et al. 2023), Nutrea (Choi et al. 2023), ToG (Sun et al. 2023), RoG (Luo et al. 2023), KAPING (Baek, Aji, and Saffari 2023), ReasoningLM (Jiang et al. 2023), FiDeLis (Sui et al. 2024), GNN-RAG (Mavromatis and Karypis 2024), DoG (Ma et al. 2025a), DualR (Liu et al. 2025) , DP (Ma et al. 2025b), and RwT (Shen et al. 2025). More introductions are provided in Appendix D.
Implementation Details. We implement the DAMR framework using PyTorch, and all experiments are conducted on NVIDIA A100 GPUs. The LLM-based planner is implemented with GPT-4.1 (Liu et al. 2023), while question and relation embeddings are generated from Qwen3-8B (Yang et al. 2025) with an embedding dimension of 1024. For the path evaluation module, we use a 128-dimensional embedding and employ the Adam optimizer with a learning rate of $1\times 10^{-4}$ during pretraining and $1\times 10^{-5}$ during fine-tuning. The model consists of two Transformer layers and is trained for 15 epochs in the pretraining stage and 10 epochs in the fine-tuning stage. Following (Luo et al. 2023; Yao et al. 2025; Ma et al. 2025b), we evaluate DAMR using Hits@1 and F1 score, assessing answer correctness and overall accuracy for questions with potentially multiple correct answers.

### v4

@@ 3.4 Context-Aware Path Evaluation
While LLM-guided expansion effectively narrows the search space by selecting semantically relevant relations, it does not ensure that all expanded paths remain valid in the broader reasoning context. As the search progresses, path semantics evolve dynamically, and early promising trajectories may later become misleading or irrelevant. Prior works attempt to filter or pre-plan paths to improve relevance (fang2024karpa, ; long2025eperm, ), but still leave gaps in capturing evolving path semantics during search. To address this, we integrate a lightweight Transformer-based path scorer into MCTS’s simulation phase, employing cross-attention to jointly encode the question and current path, enabling adaptive evaluation that reflects evolving semantics.
@@
Context-Aware Path Evaluator. In the Simulation phase, we assess the quality of candidate paths generated during MCTS rollouts. Given a question $q$ and a candidate relation path $p_{r}=(r_{1},r_{2},\ldots,r_{l})$, where $p_{r}$ is formed by sequentially selecting relations during expansion, both $q$ and $p_{r}$ are first encoded using a pre-trained LLM. Let $\mathbf{z}_{q}\in\mathbb{R}^{d}$ denote the embedding of $q$ and $\mathbf{z}_{r_{i}}\in\mathbb{R}^{d}$ the embedding of relation $r_{i}$. To capture the sequential structure of $p_{r}$, we introduce a learnable positional encoding $\mathbf{e}_{i}^{\text{pos}}$ for each relation. The final input sequence is obtained by combining $\mathbf{z}_{r_{i}}$ with $\mathbf{e}_{i}^{\text{pos}}$ and feeding the sequence into a Transformer encoder:
$\mathbf{E}_{p_{r}}=\text{Transformer}([\mathbf{z}_{r_{1}}+\mathbf{e}_{1}^{\text{pos}},\ldots,\mathbf{z}_{r_{l}}+\mathbf{e}_{l}^{\text{pos}}]),$ (3)
@@
where $\mathbf{e}_{i}^{\text{pos}}=\mathbf{E}^{\text{pos}}[i]$ is the positional embedding of the $i$-th hop, drawn from a trainable matrix $\mathbf{E}^{\text{pos}}\in\mathbb{R}^{L\times d}$, where $L$ is the maximum path length and $d$ the embedding dimension. To further incorporate question-specific information, we apply a cross-attention mechanism, allowing the encoded path representation $\mathbf{e}_{p_{r}}$ to attend to the question embedding $\mathbf{z}_{q}$:
@@
$\mathbf{H}=\mathbf{E}_{p_{r}}+\mathrm{CrossAttn}(\mathbf{E}_{p_{r}},\mathbf{z}_{q}),\;with\;\text{CrossAttn}(\mathbf{E}_{p_{r}},\mathbf{z}_{q})=\text{softmax}({\mathbf{E}_{p_{r}}\cdot\mathbf{z}_{q}^{T}}/{\sqrt{d_{k}}})\cdot\mathbf{z}_{q}.$ (4)
@@
We then apply attention pooling over the relation representations $\mathbf{H}$ to obtain the hidden state of the relation path:
$\mathbf{s}_{p_{r}}=\sum_{i=1}^{l}\alpha_{i}\textbf{h}_{i},\quad\alpha=\text{Softmax}\big(\text{MLP}(\mathbf{H})\big),$ (5)
@@
where $\mathbf{h}_{i}$ is the hidden state of the $i$-th relation and $\alpha_{i}$ its learned attention weight. This pooling allows the model to emphasize informative steps along the reasoning path. The pooled path representation $\mathbf{s}_{p_{r}}$ is then concatenated with the question embedding $\mathbf{z}_{q}$, and the combined vector is passed through a multi-layer perceptron to compute the plausibility score of the question–path pair:
$S(q,p_{r})=\mathrm{MLP}\big([\mathbf{s}_{p_{r}}:\mathbf{z}_{q}]\big).$ (6)
@@
This context-aware evaluator assigns plausibility scores to partial reasoning paths by jointly encoding the question and relation sequence, providing accurate, context-sensitive guidance for MCTS.
@@
Pre-training of Evaluator. To train the context-aware evaluator, we construct supervision by sampling positive and negative paths from local subgraphs. A path is positive if it links the head entity to a correct answer within a hop limit. Negatives come from two sources: hard negatives that terminate near but miss the answer, and random negatives from walks that avoid answer entities. Each training instance is a triplet $(q,p^{+},p^{-})$, with sequences zero-padded and masked for efficient batch training.
@@
The model computes a plausibility score $S(q,p)$ for each question-path pair and is optimized using the Pair-wise Ranking loss (rendle2012bpr, ) to encourage higher scores for positive paths:
@@
where $\sigma(\cdot)$ is the sigmoid function. This training strategy equips the evaluator with the ability to distinguish plausible reasoning paths, thereby improving the guidance signal during MCTS inference.
3.5 Path-based Dynamic Refinement
@@
To construct supervision signals for fine-tuning, we dynamically sample pseudo-path pairs ($\hat{p}^{\prime}_{i}$, $\hat{p}^{\prime}_{j}$) from the set of explored paths during MCTS. Instead of relying on the evaluator’s predictions, we assign pseudo-labels based on empirically grounded values derived from the search process. Specifically, for entity $e_{i}$ along a reasoning path $p_{r}$, we define its search value as: $w_{e_{i}}=\frac{w_{p_{r}}}{n_{e_{i}}}$, where $w_{p_{r}}$ is the cumulative reward from all rollouts passing through $p_{r}$, and $n_{e_{i}}$ is the visite count of entity $e_{i}$. Given a pair of paths, we assign pseudo-labels based on their relative values:
$\left(\hat{p}^{+},\,\hat{p}^{-}\right)=\begin{cases}(p^{\prime}_{i},\,p^{\prime}_{j}),&\text{if }w_{e_{i}}>w_{e_{j}},\\ (p^{\prime}_{j},\,p^{\prime}_{i}),&\text{otherwise}.\end{cases}$ (9)
The path evaluator is then fine-tuned using the Pair-wise Ranking loss in Eq. 7, encouraging higher scores for more promising paths.
3.6 Reasoning Process
The overall reasoning process is outlined in Appendix A. The framework first initializes a path evaluator to discriminate between plausible and implausible KG paths, providing a foundation for downstream search. During dynamic MCTS, the algorithm iteratively performs selection, expansion, simulation, and backpropagation. In expansion, an LLM-based planner adaptively selects the top-$k$ relations most relevant to the question, steering the search toward semantically meaningful paths. The evaluator guides simulation by prioritizing trajectories likely to yield correct answers, while pseudo-path pairs sampled during search are periodically used for refinement. Finally, entities reached by high-scoring paths are aggregated to form the answer set.
4 Experiments
4.1 Experimental Settings
Datasets. To evaluate the effectiveness of DAMR, we conduct experiments on two widely used KGQA benchmarks: WebQSP talmor2018web and CWQ yih2016value . Following prior work sun2023think ; liu2025dual , we uniformly sample 1,000 questions from the test sets of both datasets to evaluate the performance. More details about datasets are provided in Appendix  B.
Baselines. We compare DAMR with a comprehensive set of baselines. These baselines include: the semantic parsing methods, e.g., KV-Mem (miller2016key, ), EmbedKGQA (saxena2020improving, ), QGG (lan2020query, ), NSM (he2021improving, ), TransferNet (shi2021transfernet, ), and KGT5 (saxena2022sequence, ); the retrieval-based methods, e.g., GraftNet (sun2018open, ), PullNet (sun2019pullnet, ), SR+NSM (zhang2022subgraph, ), and SR+NSM+E2E (zhang2022subgraph, ); the general LLMs, including Flan-T5-xl (chung2024scaling, ), Alpaca-7B (taori2023stanford, ), Llama3-8B (dubey2024llama, ), Qwen2.5-7B (team2024qwen2, ), ChatGPT (schulman2022chatgpt, ), and ChatGPT+CoT (wei2022chain, ); and recent LLMs with KG methods, including UniKGQA (jiang2022unikgqa, ),DECAF (yu2022decaf, ), KD-CoT (wang2023knowledge, ), Nutrea (choi2023nutrea, ), ToG (sun2023think, ), RoG (luo2023reasoning, ), KAPING (baek2023knowledge, ), ReasoningLM (jiang2023reasoninglm, ), FiDeLis (sui2024fidelis, ), GNN-RAG (mavromatis2024gnn, ), DoG (ma2025debate, ), DualR (liu2025dual, ) , DP (ma2025deliberation, ), and RwT (shen2025reasoning, ). The details are provided in Appendix C.
@@ 4.1 Experimental Settings
Semantic Parsing KV-Mem (miller2016key, ) 46.7 34.5 18.4 15.7
EmbedKGQA (saxena2020improving, ) 66.6 - 45.9 -
QGG (lan2020query, ) 73.0 73.8 36.9 37.4
NSM (he2021improving, ) 68.7 62.8 47.6 42.4
TransferNet (shi2021transfernet, ) 71.4 - 48.6 -
KGT5 (saxena2022sequence, ) 56.1 - 36.5 -
Retrieval GraftNet (sun2018open, ) 66.4 60.4 36.8 32.7
PullNet (sun2019pullnet, ) 68.1 - 45.9 -
SR+NSM (zhang2022subgraph, ) 68.9 64.1 50.2 47.1
SR+NSM+E2E (zhang2022subgraph, ) 69.5 64.1 49.3 46.3
LLMs Flan-T5-xl chung2024scaling 31.0 - 14.7 -
Alpaca-7B taori2023stanford 51.8 - 27.4 -
Llama3-8B (dubey2024llama, ) 30.3 25.7 30.5 27.8
Qwen2.5-7B (team2024qwen2, ) 28.4 23.7 25.9 24.1
ChatGPT (schulman2022chatgpt, ) 66.8 - 39.9 -
ChatGPT+CoT (wei2022chain, ) 75.6 - 48.9 -
LLMs+KGs UniKGQA (jiang2022unikgqa, ) 77.2 72.2 51.2 49.0
DECAF yu2022decaf 82.1 78.8 70.4 -
KD-CoT wang2023knowledge 68.6 52.5 55.7 -
Nutrea (choi2023nutrea, ) 77.4 72.7 53.6 49.5
ToG (sun2023think, ) 81.9 76.0 68.5 60.2
RoG (luo2023reasoning, ) 80.8 70.8 57.8 56.2
KAPING (baek2023knowledge, ) 72.4 65.1 53.4 50.3
ReasoningLM (jiang2023reasoninglm, ) 78.5 71.0 69.0 64.0
FiDeLis (sui2024fidelis, ) 84.3 78.3 71.5 64.3
GNN-RAG (mavromatis2024gnn, ) 82.8 73.5 62.8 60.4
DoG (ma2025debate, ) 65.4 55.6 41.0 46.4
DualR (liu2025dual, ) 81.5 71.6 65.3 62.1
DP (ma2025deliberation, ) 87.5 81.4 75.8 69.4
RwT (shen2025reasoning, ) 87.0 79.7 72.4 66.7
@@
Implementation Details. We implement the DAMR framework using PyTorch, and all experiments are conducted on NVIDIA A100 GPUs. The LLM-based planner is implemented with GPT-4.1 (liu2023evaluating, ), while question and relation embeddings are generated from Qwen3-Embedding-8B (yang2025qwen3, ) with an embedding dimension of 1024. For the path evaluation module, we use a 128-dimensional embedding and employ the Adam optimizer (kingma2014adam, ) with a learning rate of $1\times 10^{-4}$ during pretraining and $1\times 10^{-5}$ during fine-tuning. The model consists of two Transformer layers and is trained for 15 epochs in the pretraining stage and 10 epochs in the fine-tuning stage. Following (luo2023reasoning, ; yao2025learning, ; ma2025deliberation, ), we evaluate DAMR using Hits@1 and F1 score, assessing answer correctness and overall accuracy for questions with potentially multiple correct answers.
4.2 Performance Comparison
We report the experimental results of DAMR in Table 1, benchmarking its performance against state-of-the-art baselines across KGQA datasets. From the results, we find that: (1) Semantic parsing and retrieval-based methods serve as early foundations for KGQA by extracting subgraphs and capturing structural semantics. However, embedding-based models struggle with complex relational patterns, while retrieval-based methods rely on rigid pipelines that limit generalization. In contrast, LLM with KG approaches combine the language understanding of LLMs with structured reasoning over KGs, enabling more flexible path exploration and improved adaptability to diverse, multi-hop queries. (2) General-purpose LLMs, such as ChatGPT and Llama3-8B, show basic reasoning ability but often perform worse than methods that combine LLMs with KGs in KGQA tasks. This is mainly because they are not grounded in domain-specific knowledge, making them more likely to produce incorrect or made-up answers. (3) DAMR consistently outperforms all baselines across both datasets, showcasing its strong reasoning capability. This superior performance is driven by its integration of an LLM-based planner, which selectively retrieves relevant relations to reduce noise and guide the search toward high-quality reasoning paths, and a path evaluation model that is dynamically fine-tuned during search to capture semantic differences among candidate paths and accurately rank those most likely to yield correct answers.
4.3 Efficiency Analysis
4.4 Ablation Study
We conduct ablation studies to assess the contributions of key components in DAMR: (1) DAMR w/o PE: removing the path evaluation module; (2) DAMR w/o FT: disabling fine-tuning of the evaluator; and (3) DAMR w/ GPT-4.1: replacing the context-aware evaluator with a general LLM.
Experimental results are summarized in Table 3. From the results, we find that: (1) Removing the path evaluation module (DAMR w/o PE) causes a noticeable performance drop on both datasets, underscoring its critical role in guiding the search process. Without this component, the model cannot effectively assess or rank candidate paths, leading to suboptimal reasoning and reduced accuracy. (2) Compared to DAMR w/o FT, the proposed DAMR consistently achieves superior results on both datasets, highlighting the importance of the finetuning mechanism in the path evaluation module. This mechanism enables the model to adapt to the evolving distribution of explored paths, improving its ability to distinguish plausible from implausible reasoning trajectories. (3) Replacing the context-aware path evaluation module with general LLMs yields degraded performance, confirming the advantage of our fine-tuned path scorer. By capturing fine-grained semantic distinctions among candidate paths, it provides more accurate evaluation signals, enhancing overall search effectiveness.